OpenStudy (lacris):

Please help me c: I will give you a super shiny medal! Let S = {1,2,3,4,... 100} with each outcome having equal probability. What is the probability that a number picked at random is at least five.

OpenStudy (lacris):

First I know the equation is a compute at least probabilities: 1-P(none) ? But the probability that none are 5 is: 100-5=95/100 = 0.95? so it is: 1-P(0.95)? What is P then?

OpenStudy (lacris):

(by none I meant the opposite of at least picking a 5)

OpenStudy (welshfella):

at leat picking a 5 means that the number picked is any number from 5 to 100 ( including 5 and 100)

OpenStudy (welshfella):

* at least

OpenStudy (holsteremission):

The sample space \(S\) follows a discrete uniform distribution, so each number \(x\) has probability \(p(x)=\dfrac{1}{100}\) of getting picked. This means the probability of selecting a number that is at least \(5\) is \[p(x\ge5)=\sum_{x=5}^{100}\frac{1}{100}=1-\sum_{x=1}^4\frac{1}{100}=1-p(x<5)\]

OpenStudy (lacris):

It was actually the classical method since all of them were equal probabilities, so: P (at least 5) = 1- P ( 4 or less) = 1-4/100 = 96/100 Thank you for helping, I asked my teacher this question and got this answer.